OpenStudy (anonymous):

If 3 times a number, added to 2, divided by the number plus 3, the result is 5/2. what is the number?

OpenStudy (anonymous):

@Jamielynne @bibby @coolman500

OpenStudy (anonymous):

Let the number be : \(n\).. 3 times a number means : \(3 \times n \implies 3n\) added to 2 means : Add that number with \(2\) : \(3n + 2\) Divided by the number plus 3 means: \[\frac{3n+2}{n+3}\]

OpenStudy (anonymous):

ok so the 3n+2/n+3 is = to 5/2?

OpenStudy (anonymous):

The result is finally 5/2. So, equate that to 5/2. \[\frac{3n+2}{n+3} = \frac{5}{2}\]

OpenStudy (anonymous):

Good.. :)

OpenStudy (anonymous):

Can you find \(n\) from here?

OpenStudy (anonymous):

um im not sure

OpenStudy (anonymous):

Okay, don't worry.. Do you know how to cross multiply?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Do that and tell me what you end up with?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so i got 15n+10=2n+6 so i need to combine like terms right?

OpenStudy (anonymous):

Oh.. No..!! Sorry..!!

OpenStudy (anonymous):

?

OpenStudy (anonymous):